Consider a spring with a spring constant of 459.8 N /m. How much work in J must be done to compress this spring a distance of 26.7 cm?
1
Expert's answer
2018-11-06T15:48:09-0500

The work done

W=(kΔx^2)/2=459.8×〖0.267〗^2/2=16.4 J